Millie and Samantha began saving money at the same time. Millie plans to save $3 a month, and Samantha plans to save $5 a month. After how many months will Samantha have saved exactly $10 more than Millie?
Answer by checkley77(12844)

86-70=16
92-86=6
6:16=3:8
86*8+92*3=86*(8+3)
688+276=86*11
964=946